OpenStudy (anonymous):

Multiply and reduce (4x^3/3)x((9)^3/(x)) A.972 B.36x² C.108

OpenStudy (austinl):

\(\dfrac{4x^3}{3}\times\dfrac{9^3}{x}=\dfrac{4x^3\times9^3}{3\times x}=~?\)

OpenStudy (anonymous):

there is parenthesis around the 9 and then it has the 3. (9)^3 im not sure if the parenthesis make a difference.

OpenStudy (austinl):

\((9)^3=9^3\)

OpenStudy (anonymous):

how do i solve it?

OpenStudy (austinl):

Like I stated above, you need to do the multiplication first, \(4x^3\times9^3=~?\) \(3\times x=~?\)

OpenStudy (anonymous):

3*x = 3x but how do i solve the other one?

OpenStudy (austinl):

What is \(9^3?\)

OpenStudy (anonymous):

729

OpenStudy (austinl):

\(729\times4=~?\)

OpenStudy (anonymous):

2916

OpenStudy (austinl):

Correctomundo, so now we have, \(\dfrac{2916\times x^3}{3\times x}\) Correct? So then we can then separate this, as such, \(\dfrac{2916}{3}\times\dfrac{x^3}{x}\) What do each of those equal?

OpenStudy (anonymous):

972

OpenStudy (austinl):

Ok, so now we have, \(972\times\dfrac{x^3}{x}\) What is \(\dfrac{x^3}{x}?\)

OpenStudy (anonymous):

im not so sure will it be just x?

OpenStudy (austinl):

\(\dfrac{x^3}{x^1}=x^{3-1}\)

OpenStudy (anonymous):

Thank you for helping me.

OpenStudy (austinl):

no problem!